• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name C. C. Chang

    Log entry time 00:09:57 on June28,2002

    Entry number 86902

    keyword=Shift Summary

    17:05 Left arm HV alarm came on. We determined it was just the readout
    since all rates remain normal. Called Javier (and Jack). Jack told
    us what to do over the phone. After pushing Arcnet reset button
    all HV displays went blanck. the display should have come back;
    it hasn't. Jack is looking into the problem. Meanwhile, we continue
    to taking data since (we believe) the HVs are still on (det. replay
    looks good).
    17:25 Just notice that left arm Q1 tripped.
    17:45 Left arm Q1 is back.
    17:50 Request hall access to reset left arm HV crate and HV IOC
    located in right arm hut to solve problem in HV display.
    19:00 It turns out the problem was more complicated than we thought.
    We were unable to restore the connection for HV display
    (see HALOG entry # 86890 for more detailed explanation). We are
    done in the hall. Called MCC to restore beam.
    19:10 MCC called to give us CW beam.
    19:20 Resume data taking for Q2_l50.
    20:30 The software person who were here to fix the HV display problem
    earlier called to inform us that the HV display is back (without
    him doing anything, strange!)
    22:30 Continue to take data for Q2_l50.
    23:18 ROC14 hanged. Reboot ROC14.
    24:00 Finish Q2_l50.